# Modelling reliability under deep decarbonisation of the European electricity grid

### v3

@@ Abstract
We introduce a new model able to characterise the relative costs of low-carbon generation mixes when subject to varying grid-reliability targets, which are evaluated using 43 years of historical weather data. The model can be used to develop cost-minimising strategies, and to assess the extent, duration, and frequency of outages under different technology mixes. For a 35-country subset of the European electricity grid, we show that reliably meeting the last 1% of demand accounts for about 36% of the entire system cost when only onshore wind, solar, and storage are used in the generation mix. We also show that by including small amounts of low-cost, dispatchable load-following generation (e.g., natural gas, hydro, or similar) the cost of a reliable, high-renewables grid can be drastically reduced. For example, the system cost can be reduced by 31% when just 1% of generation is allowed to come from natural gas.
keywords:
Long-term grid reliability , Energy system modelling , Optimising renewable generation mix , Net Zero in Europe , Decarbonisation trade-offs
{highlights}
First multi-decadal simulation of a renewable European electricity grid
Finds that both storage and dispatchable backup are crucial for maintaining grid reliability
Dispatchable backup can enable large cost-savings at low utilisation rates
Cost-optimised scenarios are dominated by wind rather than solar
1 Introduction
@@
Because of their low and falling costs, wind and solar power are expected to play a major role in the decarbonisation of electricity systems worldwide [1, 2, 3]. However, the inherent variability of wind and solar generation has led to questions about their potential to achieve very deep decarbonisation (near 100%) while also maintaining grid reliability at reasonable cost [4, 5, 6, 7, 8]. These concerns have motivated several studies on the feasibility of a 100% renewable power sector [9, 10, 11, 12, 13] and energy sector [3, 14, 15, 16] in Europe. Such studies generally estimate wind and solar generation potential across the European continent using climate reanalysis datasets, and feed this into a power system model to optimise for the minimum-cost system that meets demand given technical constraints on generator capacity, transmission capacity, generator ramp-time and other factors. However, these models are computationally expensive and, as a result, are often limited in their temporal scope. For example, two recent and widely-cited studies exploring 100% renewable scenarios in Europe employed the LUT energy system model [16] and the PLEXOS model energy system model [13]. Both studies were conducted using only a single year of weather data: the LUT study used a typical weather year [17], while the PLEXOS study used a performance-limiting weather year but included just one week of data per month during the year.
@@
This reduction in temporal scope is necessary when using complex power sector models like PLEXOS and LUT, but it means that the results obtained are not robust over multi-decadal time scales. The need for high temporal resolution, long-duration weather data in energy system planning is already well-documented [18, 19, 20], and recent work by Ruggles et. al. has found that almost 40 years of weather data are required to ensure a reliable grid in systems without dispatchable backup [20]. Even when dispatchable generation is allowed to supply 5% of all electricity, 15 years of weather data are still required to ensure a reliable grid.
@@
Computationally intensive models force other limitations as well. The LUT model assumes perfect foresight of supply and demand [17] in its simulations, which yields overly optimistic projections surrounding the levels of storage and backup needed to supplement a highly renewable grid. In addition, detailed sensitivity analyses are often impossible with computationally expensive models. While previous studies have explored the sensitivity of the system cost to single parameters like the permitted carbon intensity of the grid or individual technology costs [5, 21], scanning across multiple inputs simultaneously (e.g., investigating how varying the carbon intensity would affect the overall system cost under many different technology cost assumptions) becomes prohibitively expensive.
@@
In this paper, we use RESCORE to give the first cost estimate of a highly-renewable grid in Europe with a realistic reliability constraint over a multi-decadal time period. Modern targets for grid reliability in Europe are 99.97% [23], but since desired grid reliability is an input to the RESCORE model, we are also able to explore for the first time how varying this reliability target would affect both the cost and make-up of a highly-renewable European grid. Furthermore, RESCORE returns full reliability information including the severity and duration of all outages, allowing these metrics to be analysed in detail. The speed of the model also allows for a rich sensitivity analyses in which the interplay between several different constraints and inputs can be investigated, which helps find global optimums faster. In this paper, we investigate the effect of simultaneously varying the target grid reliability, the carbon intensity, and the cost of electricity storage on the overall system cost. These constraints are highly interdependent, and must be studied together using a tool like RESCORE in order to fully capture their impact on the system cost.
Figure 1: Map of European countries included in the analysis. Countries shaded green are included in both our main analysis (presented in this paper) and in a validation analysis performed using the renewables.ninja dataset (see Appendix A and Supplementary Information) [24, 25]. Countries shaded blue are included in our primary analysis but not in the renewables.ninja analysis. Countries shaded orange are connected to the Central European Synchronous Area (CESA) but are not included in either of our analyses.
2 Methods I - Data
@@
2.1 Calculating solar generating potential

### v4

@@ Abstract
Conventional feasibility studies of deep decarbonisation are often limited in their temporal scope, and are thus unable to draw conclusions about grid reliability over multi-decadal time periods. To address this problem, we introduce RESCORE, a fast and transparent model that uses 43 years of hourly weather data to evaluate both the cost and reliability characteristics of low-carbon electricity mixes. Applying RESCORE to the European electricity grid, we show that infrequent but extreme weather events play an outsized role in setting the cost for generation mixes without dispatchable backup: reliably meeting the last 1% of demand accounts for about 36% of the entire system cost when the mix contains only solar, onshore wind and storage. We also show that by including small amounts of low-cost, dispatchable generation (e.g., natural gas, hydro, or similar) the cost of a reliable, high-renewables grid can be drastically reduced.
I Introduction
Europe has some of the most aggressive CO2-reduction targets in the world, with the EU pursuing net zero carbon emissions by 2050 across its 27 member states  [1] and several individual countries setting legally binding targets to reach the same milestone [2, 3, 4]. Due to their low and falling costs, wind and solar power are expected to play a major role in the decarbonisation of the electricity grid in Europe [5, 6, 7].
@@ I Introduction
However, the inherent variability of wind and solar generation has led to questions about their potential to achieve very deep decarbonisation (near 100%) while ensuring reliable grid operation [8, 9, 10, 11, 12]. These concerns have motivated several studies on the feasibility of a 100% renewable power sector [13, 14, 15, 16, 17] and energy sector [7, 18, 19, 20] in Europe. Such studies generally estimate wind and solar generation potential across the European continent using climate reanalysis datasets, and feed this into a power system model to optimise for the minimum-cost system that meets demand given technical constraints on generator capacity, transmission capacity, generator ramp-time and other factors. However, these models are computationally expensive and, as a result, are often limited in their temporal scope. For example, two widely-cited studies exploring 100% renewable scenarios in Europe employed the LUT energy system model [20] and the PLEXOS model energy system model [17]. Both studies were conducted using only a single year of weather data: the LUT study used a typical weather year [21], while the PLEXOS study used a performance-limiting weather year but included just one week of data per month during the year.
@@
This reduction in temporal scope is necessary when using complex power sector models like PLEXOS and LUT, but it means that the results obtained are not robust over multi-decadal time scales. The need for high temporal resolution, long-duration weather data in energy system planning is already well-documented [22, 23, 24], and recent work by Ruggles et. al. has found that almost 40 years of weather data are required to ensure a reliable grid in systems without dispatchable backup [24].
@@
Computationally expensive models force other limitations as well. The LUT model assumes perfect foresight of supply and demand in its simulations [21], which yields overly optimistic projections surrounding the levels of storage and backup needed to supplement a highly renewable grid. In addition, detailed sensitivity analyses are often impossible with computationally expensive models. While previous studies have explored the sensitivity of the system cost to single parameters like the permitted carbon intensity of the grid or individual technology costs [9, 25], scanning across multiple inputs simultaneously (e.g., investigating how varying the carbon intensity would affect the overall system cost under many different technology cost assumptions) becomes prohibitively expensive.
@@
In this paper, we use RESCORE to give the first cost estimate of a highly-renewable grid in Europe with a realistic reliability constraint over a multi-decadal time period. Modern targets for grid reliability in Europe are 99.97% (corresponding to an average of 3 hours per year in which supply cannot meet demand) [27], but since desired grid reliability is an input to the RESCORE model, we are also able to explore for the first time how varying this reliability target would affect both the cost and make-up of a highly-renewable European grid. RESCORE returns full reliability information including the severity and duration of all outages, enabling these metrics to be analysed in detail. The speed of the model also allows for a rich sensitivity analysis in which the interplay between several different constraints and inputs can be investigated. We investigate here the effect of simultaneously varying the target grid reliability, the carbon intensity, and the cost of electricity storage on the overall system cost. These constraints are highly interdependent, and must be studied together using a tool like RESCORE in order to fully capture their impact on the system cost.
II Methods I - Data
@@
II.1 Calculating wind & solar generating potential
An estimate for the wind and solar generating potential at every grid was obtained using data from the MERRA-2 Radiation Diagnostics [28] and Single-Level Diagnostics [29] datasets. For solar generation, we assumed non-tracking panels and a standard efficiency of 21% (defined as the efficiency under 1000 $W/m^{2}$ direct illumination at a panel temperature of 25 °C) [31]. To calculate the generation, we followed a similar method to Pfenninger and Staffell [32], and Sandia National Laboratory [33]—and we accounted for the temperature-dependence of panel efficiency using Huld’s model [34]. Full details of the calculation can be found in Appendix A.
A similar calculation was performed for the wind generating potential. We assumed all wind turbines to have a turbine rotor diameter of 110 m, a hub-height of 100 m, a maximum rated power of 4.1 MW and cut-in and cut-out speeds of 3 m/s and 25 m/s respectively. Hellman’s power law equation was used to extrapolate wind speeds recorded at 10m and 50m in the MERRA-2 dataset to the assumed hub height of 100 m [35]. Again, full details can be found in Appendix A.
This process yielded two hourly-resolution generation profiles at every MERRA-2 grid point: a ‘wind power per turbine’ and a ‘solar power per unit area’. We averaged over all 2044 points in the grid to obtain a single solar generation profile and a single wind generation profile for the whole of Europe, and then scaled these profiles to set the installed capacity of wind and solar in the baseline case equal to peak electricity demand in Europe (486 GW).
II.2 Estimating electricity demand
The European Network of Transmission System Operators for Electricity (ENTSO-E), which operates the synchronous grid in Europe, requires its members to publish hourly electricity demand data every year, and makes this information publicly available [30]. However, not all countries connected to the synchronous European grid (CESA) are members of ENTSO-E, and the UK, Turkey and the North African states in particular do not publish hourly demand data on the ENTSO-E platform. Data for the UK are available elsewhere [36], but Turkey and the North African states were excluded from our analysis due to data being unavailable. Some countries, such as Ukraine and Albania, use the ENTSO-E platform but do not reliably publish data. We only included in our analysis the 35 countries that provided demand data for more than 8700 hours of the year (out of 8760) in 2022—see Appendix C. The borders of these countries define the locations of the MERRA-2 grid points used to calculate solar and wind generating potential.
II.3 Storage & Backup
In order to analyse the effects of adding storage and dispatchable backup to a highly renewable electricity grid, we had to choose technologies to capture these characteristics. For storage, we chose lithium-ion batteries, since this is currently the lowest cost battery technology and has been widely studied in the context of future grid-scale electricity storage [37, 38]. We followed the cost estimates in [39] for a 4-hour Li-ion storage system, and assumed that maintenance covered under the fixed annual O&M costs ensures no degradation of the battery over its life cycle [39]. For simplicity, we have assumed perfect round-trip efficiency and no self-discharge for the batteries. We would not expect any significant changes to our results if these terms were included since, in our threshold-based model (see Section III.1), storage inefficiencies would lead to dispatchable backup being deployed slightly earlier. This would increase fossil fuel emissions for each scenario slightly, but would likely not affect the system cost or the reliability of the grid.
III Methods II - The RESCORE Model
The result of the procedure detailed in Section II is a dataset containing average solar generation, average wind generation and average electricity demand at hourly resolution over the 43-year period. We feed this dataset as an input to RESCORE, which uses a rules-based dispatch model to calculate the average grid reliability and the percentage of electricity produced by the dispatchable source for a given grid configuration. We then combine this with a cost model for system to find the minimum-cost blend given constraints on grid reliability and dispatchable generation deployment. While not essential to the model, including the constraint on dispatchable deployment allows us to explore the cost and reliability trade-offs associated with different levels of decarbonisation if natural gas is used as the dispatchable generator. We do not include transmission in our cost estimates, since we do not model transmission, but achieving the perfect levels of interconnectivity assumed here would require major upgrades to the European transmission grid.
III.1 Running the simulation
  1. 1.
    The renewables overbuild factor
  2. 2.
    The relative proportion of wind and solar in the generation mix
  3. 3.
    The total system storage capacity in GWh
  4. 4.
    The total natural dispatchable generation capacity in GW
  5. 5.
    A threshold to indicate when dispatchable generation should start being used (expressed as a percentage of the total system storage capacity)
At each hour, we calculate solar and wind generation based on the overbuild factor (where a $\times$1 overbuild factor means the installed capacity for wind and solar equals peak demand) and the respective proportions of wind and solar in the generation mix. If residual load (demand minus renewable generation) is less than zero, then the excess renewable generation is used to replenish the storage up to its maximum value. If residual load is greater than zero and storage levels are greater than the threshold, then storage is called first to help meet demand. If residual load is greater than zero and storage levels are below the threshold, then the dispatchable generators are turned on to meet demand and to restore the storage levels back to the threshold.
The simulation cycles through each hour in the 43-year period and outputs the reliability (the percentage of hours in which supply was sufficient to meet demand), as well as the percentage of total electricity that was generated by dispatchable generation.
1) Never use dispatchable backup unless storage levels are below the threshold.
III.2 System cost calculations
As well as running the grid simulation described in Section III.1 for each set of generation capacity inputs, the corresponding annual system cost (neglecting transmission costs) is also calculated.
Solar 790 [42] 10 [43] 25 [42]
Wind 1540 [42] 40 [43] 25 [42]
Storage 200 (\$/kWh) [39] 10 (\$/kWh-year) [39] 15 [39]
Dispatchable backup (natural gas) 1000 [42] 20 [44] 30 [45]
Table 1: Cost estimates for each technology used in the analysis. Sources are IEA [42], IRENA [43], EIA [44] and NREL [39, 45]. Variable Operations & Maintenance (O&M) costs were assumed negligible for all technologies, and fuel costs for natural gas were neglected because its maximum utilisation was just 2% of all electricity generation in the study. All cost data in $\mathdollar_{2025}$. Figure 5 shows a sensitivity analysis to storage costs. The conversion to investment cost is described in the Methods section.
$A=P_{0}\left(\dfrac{e^{rL}{(e^{r}-1)}}{e^{rL}-1}\right),$ (2)
III.3 Limitations of the RESCORE model
Figure 1: Map of European countries included in the analysis. Countries shaded green are included in both our main analysis (presented in this paper) and in a validation analysis performed using the renewables.ninja dataset (see Appendix B and Supplementary Information) [46, 32]. Countries shaded blue are included in our primary analysis but not in the renewables.ninja analysis. Countries shaded orange are connected to the Central European Synchronous Area (CESA) but are not included in either of our analyses.
Finally, there is an extensive discussion in the literature surrounding the validity of reanalysis datasets like MERRA-2 for renewable generation studies [46, 32, 47, 48]. Staffell and Pfenninger found discrepancies between wind and solar generation estimates from MERRA-2 and real-world generation data for European countries [46, 32]. To validate the conclusions of this paper, we have performed our analysis using both uncorrected MERRA-2 data as well as renormalised datasets for wind and solar published by Staffell and Pfenninger on the renewables.ninja platform [46, 32]. The countries covered in our original analysis and the countries covered in our validation analysis with the renewables.ninja dataset are shown in Figure 1. Full details of the differences between the datasets are included in Appendix B, and a direct comparison of the results between the two datasets is included in the Supplementary Information. The results and conclusions presented in this paper were found to be robust to the choice of dataset used.
IV Results
IV.1 Required Overbuild & Storage to Meet Reliability Targets
[…]
Adding some energy storage (but no dispatchable generation) into the mix rapidly reduces the overbuild factors needed to reach target levels of grid reliability. Figure 3b plots the renewable overbuild factor required to meet different reliability targets as a function of the storage capacity in the system. For a 99.97% reliability target, 1458 GWh of storage (equivalent to 3 hours at peak demand) is sufficient to reduce the required overbuild factor from $\times$10.6 to $\times$7. However, once storage capacity extends beyond a few hours of peak demand, it has less of an effect on the required overbuild factor. A full 24 hours of storage (11,664 GWh) only reduces the required overbuild factor for 99.97% reliability to $\times$5.
We also observe in the high (>>99.9%) reliability scenarios that reliability issues are generally caused by a small number of extreme events: in the 99.97% reliable case with 24 hours of energy storage, only 3 years out of 43 see any outages at all. This highlights the need for multi-decadal analyses when considering the reliability of highly-renewable electricity grids.
IV.2 Cost-optimised scenarios
Figure 3 makes clear that some form of dispatchable generation will likely be needed to avoid prohibitively high renewable overbuild factors in a wind/solar dominated grid. In this section, we add dispatchable backup into the simulations in the form of natural gas peaking plants, and find the cost-optimised mix for solar, wind, gas and storage under different constraints on carbon intensity, grid reliability, and storage cost.
As described in the Methods section, we perform the cost-optimisation not just over generation and storage capacity, but also over the natural gas deployment threshold (see Section III.1). Essentially, this threshold parameter sets the balance between reliability and carbon-constraints. A higher threshold will reduce the risk of blackouts, while a lower threshold will reduce the amount of electricity that is generated by natural gas and thus reduce the carbon intensity of the grid.
The model is sufficiently cheap to run that we can scan over all five input parameters with high resolution. After this, each set of input parameters can be labelled with an associated cost, average reliability over the 43-year period and percentage of electricity generated by dispatchable backup. We then choose the optimum system configuration as the lowest-cost set of five inputs that satisfies given constraints on the reliability and backup generator utilisation.
IV.2.1 System costs vs. grid reliability
Figure 4: a) Optimised annual system cost versus permitted Loss of Load Expectation per year. Loss of Load Expectation is defined as the average number of hours per year in which total demand exceeds total supply. Each subplot shows specific details about a given scenario: b) 0% of generation from natural gas at 99.97% grid reliability; c) 0% from natural gas at 99% grid reliability; d) 1% from natural gas at 99.97% grid reliability; e) 1% from natural gas at 99% grid reliability. For each scenario, i) shows the cost-breakdown of the generation mix, ii) shows the percentage of total demand met during each outage (a measure of the fraction of the load that can still be served through rolling blackouts or, inversely, a measure of the demand curtailment required to avoid the outage), and iii) shows the histogram of outage durations over the 43-year simulation period.
The blue line in Figure 4a indicates that, in a European energy system containing only wind, solar and storage, the cost difference between a 99% reliable grid and a 100% reliable grid is about \$200 billion annually—or 36% of the entire system cost. As soon as dispatchable backup is added into the mix, not only are major cost savings achieved, but the cost savings associated with relaxing the grid reliability target are also reduced. When just 1% of total electricity generation is permitted to come from the dispatchable source, the optimal cost of a 99.97% reliable grid falls from \$491 billion annually to \$335 billion annually. Furthermore, with 1% of electricity from dispatchable backup, this 99.97% reliable grid can be achieved at relatively little additional cost compared with a 99% reliable grid (an additional \$13 billion per year).
These conclusions are consistent with the results in [9], which found that the costs of decarbonising the power sector in the absence of firm resources grow rapidly for the last few percent of demand.
From the cost breakdowns in Figure 4di and Figure 4ei, it is clear that the \$13 billion cost of increasing the reliability from 99% to 99.97% is the result of a doubling of natural gas capacity (not usage)—from 121 GW to 218 GW—while the rest of the infrastructure required in the cost-optimised case remains largely the same. When dispatchable backup is allowed into the system, the installed backup capacity becomes the primary lever for increasing grid reliability. The same picture emerges when we examine the 2% natural gas case in Figure 4a, with the \$10 billion difference in cost (\$299 billion annually to \$309 billion annually) between a 99% reliable grid and a 99.97% reliable grid being solely due to a doubling in the dispatchable backup capacity.
It is also interesting to note that wind significantly dominates the overall cost in all scenarios, and that all scenarios rely on large renewable overbuild factors ($\times$7 for Figure 4b, $\times$5.8 for Figure 4c, $\times$5.2 for Figure 4d and $\times$5 for Figure 4e, with overbuild defined as nameplate capacity divided by peak demand), even when storage and dispatchable backup are permitted in the system. The high share of wind generation is likely due to a combination of two factors. The first is that European electricity demand is about 30% higher in winter than in summer (as shown in Figure 2b), meaning that it aligns with the seasonal peak in wind generation rather than solar generation. The second reason is that wind generation experiences less variation than solar, with generation during the peak months approximately $\times$2 higher than during the trough (as opposed to approximately $\times$5 for solar). It makes sense, then, that a wind-dominated generation mix is most able to track the demand curve in Figure 2b with the minimum amount of overbuild. This optimum mix could change in the future if increased air conditioning use leads to a shift in peak demand from winter to summer.
Our results also show that the addition of some form of dispatchable backup into the mix affects the nature of the system outages. Comparing like-reliability cases (4b with 4d and 4c with 4e), it is clear that for cases in which 1% of electricity is allowed to come from the dispatchable source, the typical duration of an outage is shorter and the percentage of demand met during each outage is higher. Indeed, it is possible that many of the outages in which greater than 90% of demand is met could be effectively managed in the future with suitable demand-response measures.
IV.2.2 System costs vs. storage capital costs
These results show that the cost reductions arising from the inclusion of a firm and flexible backup generator significantly outweigh the cost reductions from any decrease in storage cost. For a grid with no dispatchable backup, reducing the capital cost of storage from \$400 per kWh down to \$10 per kWh results in annual system cost savings of \$93 billion (\$515 billion to \$422 billion). If, instead, the capital cost of storage is kept fixed at \$400 per kWh but natural gas is allowed to generate 1% of total electricity, then annual system cost savings of \$168 billion are achieved (\$515 billion to \$347 billion). Indeed the cost savings that result from a further 1% increase in gas generation are comparable to the cost savings from reducing the storage capital costs from \$400 per kWh to \$10 /kWh again. A grid with 1% of electricity from natural gas and storage at \$10 per kWh costs \$311 billion, while a grid with 2% of electricity from natural gas but storage at \$400 per kWh costs just \$318 billion—only a \$7 billion annual increase in the total system cost despite a 40-fold increase in the storage capital cost.
Furthermore, just as Figure 4 shows that adding dispatchable backup makes the system cost less sensitive to the target grid reliability, Figure 5 shows that adding dispatchable backup makes the system cost less sensitive to the capital cost of storage. Without backup, increasing storage capital costs from \$10 per kWh to \$400 per kWh increases system costs by 22%, compared to a 12% increase when gas is allowed to generate 1% of all electricity.
An important point to note is that non-negligible storage capacity (amounting to a couple of hours of peak demand) is still included in the cost-optimised scenario at \$400 per kWh, even when dispatchable backup is permitted. This suggests that some storage is crucial to a low-cost system and supports the conclusions of [40], which found that adding small amounts of storage to the grid dramatically reduces the energy that must be covered by dispatchable backup resources. If a direct, carbon-free substitute for gas can be obtained, then the role for storage to play will likely be very small. However, if natural gas remains the backup generator of choice, then storage will be crucial to help limit the use of natural gas to cases where it is absolutely necessary.
The patterns of outages under varying storage costs are similar to those seen in Figure 4. For cases 5d and 5e with dispatchable backup permitted, there are no outages over the entire 43-year period in which less than 70% of demand is met. Without backup, however, cases 5b and 5c show that hours where less than 70% of demand is met do occur. Furthermore, Figure 5biii shows one outage in which supply falls short of demand for over 24 hours continuously. When dispatchable backup is added to the system, no outage lasts longer than 10 hours.
Counter-intuitively, in cases with equal dispatchable backup constraints (5b/5c and 5d/5e), reducing the cost of storage causes the intensity of outages to worsen. The case shown in 5c (high storage cost) has both a higher proportion of demand met during outages and fewer long-duration outages relative to 5b (low storage cost). This is because more wind and solar capacity is deployed in the high-storage-cost case to compensate for the reduced storage capacity. When storage levels fall to zero and an outage occurs, the greater wind/solar in the high-storage-cost case means that a greater proportion of demand can be met.
V Conclusions
We have simulated a renewables-dominated European electricity grid using the RESCORE model: a new long-duration, high temporal resolution grid model, and have found that a reliable grid will require high levels of renewable overbuild—with installed peak generating capacity greater than four times the peak demand. Even so, we find that cost-minimised systems require both storage and dispatchable backup generation to achieve reliability targets. A large quantity of one cannot compensate for the complete absence of the other without incurring extreme penalties in system costs.
While backup generation could be provided by a variety of technologies, natural gas is likely to be the preferred choice in the short-to-medium term due to its low capital cost. Our results show that natural gas backup can yield large (>>30%) reductions in the system cost even at extremely low utilisation rates (producing just 1–2% of all electricity).
Appendix A Full details of Solar & Wind generation calculations
A.1 Solar